A culture of yeast doubles in size every 20 minutes. The size of the culture is now N0. Find its size in 12 hours.
Answer by stanbon(48558)

A culture of yeast doubles in size every 20 minutes. The size of the culture is now No. Find its size in 12 hours.
------------
Let t be number of hours.
A(t) = No*(2)^(60t/20)
A(t) = No*2^(3t)
---------
A(12) = No*2^(36)
A(12) = 6.8719...x 10^10 No
-----------------------------------

Cheers,
Stan H.